Solve the equation by using the square root property.
3x^2 -5 = 0
What is the solution set?
Can someone please help I don't understand the square root property? Thank you

Answer by stanbon(75887)

3x^2 -5 = 0
----
3x^2 = 5
3x^2 -5 = 0
---
x^2 = 5/3
---
x = +/- (1/3)sqrt(15)
================================
Cheers,
Stan H.
===========